# Rethinking Graph Structure Learning in the Era of LLMs

### v3

@@ 4.4 Leaf-oriented Two-step Sampling
Topology Uncertainty Sampling. As we all know, the core of GSL is to eliminate graph structural noise to enhance data utility. Notably, this noise typically does not pervade the entire graph but exists in certain graph substructures. According to Theorem 4, nodes with higher structural entropy are more likely to reside in topologically uncertain regions and thus benefit more from supervision-aware structural refinement. Therefore, to improve running efficiency, we leverage structural entropy to quantify the priority of nodes in need of structural optimization. Specifically, within each leaf community $\mathcal{C}^{\ell}(i)$, we use structural entropy $\mathcal{H}^{\mathcal{T}}(\mathcal{G},\alpha)$ to assign probabilities to each leaf node for node-wise sampling:
$\mathcal{P}_{topo}(\alpha)=\frac{\mathrm{exp}(\mathcal{H}^{\mathcal{\mathcal{T}}}(\mathcal{G},\alpha))}{{\underset{\gamma\in\mathcal{C}^{\ell}(\alpha)}{\sum}}{\mathrm{exp}(H^{\mathcal{\mathcal{T}}}(\mathcal{G},\gamma))}}.$ (6)
Semantic Similarity Sampling. To ensure practical performance, we filter the remaining leaf nodes (excluding $\alpha$) by LLM-empowered semantic similarity $y^{cls}$ to obtain a $\theta$-sized candidate set. The node $\beta$ in this candidate set is utilized for edge addition or removal with $\alpha$. For edge addition, the sampling probabilities are ranked in descending order of similarity, while for edge removal, they are ranked in ascending order. The above edge-wise sampling is defined as:
$\mathcal{P}_{sema}^{\alpha}(\beta)=\frac{\exp(\operatorname{sim}(y^{cls}_{\beta},{y}^{cls}_{\alpha}))}{{\underset{\gamma\in\mathcal{C}^{\ell}(\alpha),\gamma\neq\alpha}{\sum}}\!\!\!{\mathrm{exp}(\mathrm{sim}(y^{cls}_{\gamma},y^{cls}_{\alpha}))}}.$ (7)
@@ 5.1 Experiment Setup
We evaluate LLaTA and 13 GSL baselines on 10 widely adopted TAG datasets across multiple domains. The details on these datasets and baselines can be found in Appendix G-H. Due to space limitations, the hyperparameter settings used in the experiments are provided in the Appendix I.
@@ 5.2 Performance Comparison
Table 1: Node classification accuracy(%) on 10 TAG datasets. Shown is the $mean\pm std$ of 10 runs with different random seeds. Highlighted are the top first, second, and third accuracy. OOM denotes out of memory and OOT denotes time limits of 72 hours exceeded.
Method Cora Citeseer Pubmed WikiCS Instagram Reddit Ratings Child History Photo IDGL 86.81±0.31 75.39±0.27 87.25±0.22 79.78±0.14 63.68±0.09 65.03±0.13 44.01±0.29 45.33±0.26 80.46±0.31 82.33±0.33 SLAPS 79.89±0.50 73.51±0.64 86.41±0.49 72.50±0.27 60.91±0.13 OOM 40.76±0.23 47.46±0.21 OOM OOM GAUGO 85.24±0.27 75.24±0.43 87.12±0.38 70.74±0.29 63.34±0.21 OOM 40.19±0.24 47.27±0.19 OOM OOM HESGSL 85.91±0.45 76.03±0.31 87.17±0.27 73.29±0.22 63.83±0.16 OOM 38.71±0.42 47.81±0.25 OOM OOM SEGSL 86.90±0.54 76.52±0.20 87.42±0.38 79.68±0.29 65.26±0.21 64.87±0.26 43.20±0.47 51.80±0.39 OOT OOT ProGNN 84.18±0.23 75.08±0.21 86.89±0.14 71.90±0.16 64.45±0.19 OOM OOM OOM OOM OOM SUBLIME 84.81±0.33 75.55±0.47 87.63±0.70 78.06±0.38 64.78±0.24 58.82±0.21 39.90±0.17 49.54±0.16 OOM OOM STABLE 84.21±0.43 75.34±0.60 87.27±0.39 76.93±0.33 65.66±0.18 59.78±0.24 40.55±0.10 50.90±0.09 OOM OOM CoGSL 86.16±0.45 75.45±0.36 OOM OOM OOM OOM OOM OOM OOM OOM BORF 87.08±0.18 76.43±0.17 87.45±0.11 80.56±0.08 64.07±0.06 63.07±0.09 42.72±0.60 50.92±0.54 82.83±0.14 82.63±0.15 DHGR 86.72±0.42 76.18±0.14 86.76±0.33 79.25±0.17 64.26±0.26 66.41±0.29 43.33±0.36 52.41±0.40 82.61±0.16 81.92±0.18 GraphEdit 86.26±1.06 76.33±1.12 87.14±0.29 79.92±0.63 64.23±1.01 OOT OOT OOT OOT OOT LLM4RGNN 86.73±0.73 76.96±0.58 87.37±0.21 OOT OOT OOT OOT OOT OOT OOT LLaTA (Ours) 88.26±0.26 78.21±0.18 88.39±0.23 81.58±0.20 66.73±0.13 67.60±0.19 44.47±0.08 62.28±0.56 85.28±0.24 85.41±0.24
@@
Method Cora Citeseer Pubmed Instagram GCN 66.42±0.16 70.28±0.14 78.95±0.07 64.59±0.01 SEGSL 67.70±0.17 71.15±0.12 79.05±0.04 64.63±0.01 BORF 67.54±0.16 70.62±0.10 79.46±0.03 64.49±0.02 DHGR 66.85±0.63 70.14±0.21 78.64±0.15 64.52±0.06 LLM4RGNN 67.63±0.11 71.72±0.13 78.90±0.06 OOT LLaTA 68.39±0.14 72.18±0.15 80.25±0.06 65.37±0.02
@@ 5.4 Robustness Analysis
Downstream Backbones. According to Table 5, LLaTA consistently outperforms the base models across all datasets, achieving notable improvements of up to 10.18% for GCN and 9.31% for GAT. Although the gains for GraphSAGE are smaller, this can be attributed to its strong inherent node representation capability. Overall, these results demonstrate LLaTA’s robustness in enhancing various GNN backbones, highlighting its broad applicability across mainstream architectures.
Table 4: Improvements (compare to GNN) of LLaTA with different GNN backbones.
Dataset LLaTAGCN LLaTAGAT LLaTASAGE Reddit 67.60±0.11 63.80±0.13 60.73±0.13 Child 62.28±0.12 62.65±0.12 63.53±0.14 History 85.28±0.12 85.30±0.12 84.67±0.11 Photo 85.41±0.13 85.38±0.12 82.20±0.12 Improvement $\uparrow$ 1.79$\sim$10.18 $\uparrow$ 1.38$\sim$9.31 $\uparrow$ 0.56$\sim$9.35
@@
Method Time complexity SEGSL $\mathcal{O}(n^{2}+nlog^{2}n+n)$ GraphEdit $\mathcal{O}(pt_{\text{LLM-r}}+qd^{2}+n^{2}+(m+nk)t_{\text{LLM-i}})$ LLM4RGNN $\mathcal{O}(pt_{\text{LLM-r}}+mt_{\text{LLM-i}}+qd^{2}+n^{2})$ LLaTA $O(nlog^{2}n+n|\mathcal{C}|^{2}+nt_{\text{LLM-i}}+n)$
Figure 3: Node classification performance on the Cora and Citeseer dataset under real-world scenarios (sparse [Edge Removal] and noise [Edge Addition]).
Real-world Scenarios. To evaluate the robustness of LLaTA against sparsity and noise scenarios in practical applications, we randomly remove or add edges to the original graph structure. As illustrated in Fig. 3, LLaTA demonstrates strong predictive performance in both scenarios. In the sparse setting, LLaTA achieves optimal performance across most perturbation levels. In the noisy scenario, LLaTA performs competitively, with only LLM4RGNN surpassing it, as the latter is specifically designed to handle graph adversarial attack settings. Notably, as the edge addition rate increases, the performance of LLaTA gradually deteriorates, which can be attributed to the irreversible negative impact of such perturbations on the tree initialization.
Hyperparameter. We conducted a comprehensive analysis of the four key hyperparameters of LLaTA: $K,\epsilon,\theta$, and $r$. The experimental results for $K,\theta$, and $r$ are presented in Fig. 5, while the analysis of $\epsilon$ and the experimental results of $\theta,r$ on other datasets are provided in Appendix L. In Fig. 5 (left), we see that the optimal encoding tree height $K$ varies by dataset, with $K=3$ for WikiCS and Pubmed, and $K=2$ for Instagram. However, when $K$ increases to 5, performance significantly drops across all datasets, suggesting that excessive tree depth may cause overfitting or loss of structural information. In Fig. 5 (right), we analyze LLaTA’s sensitivity to the candidate set size $\theta$ in semantic similarity sampling and the two-step sampling frequency $r$ on the Cora dataset. The results indicate that an $r$ value between 3 and 10 yields higher accuracy, corresponding to an appropriately improved graph density. In contrast, LLaTA is less sensitive to the choice of $\theta$, achieving stable performance within the range of $\theta\in[5,15]$. A smaller $\theta$ may result in insufficient semantic similarity sampling, whereas a larger $\theta$ could introduce noise into the candidate set, negatively impacting performance.
Figure 4: Hyperparameter analysis of $K$, $\theta$ and $r$. $\theta$ and $r$ are analyzed on the Cora dataset.
@@ 5.5 Efficiency analysis
In Table 5, we analyze the theoretical time complexity of LLaTA and existing LLM-based GSL methods. Where $q$ and $p$ is the number of samples for fine-tuning and edge predictor training, respectively. $t_{\text{LLM-r}}$ and $t_{\text{LLM-i}}$ denote the fine-tuning and inference time of LLM, respectively. The magnitude of $t_{\text{LLM-i}}$ should be adaptively determined according to the specific requirements of different inference tasks. For LLaTA, The time complexity can be divided into three parts: (a) The complexity for tree construction in Sec. 4.2 is $\mathcal{O}(nlog^{2}n)$, (b) The complexity of tree-prompted LLM inference in Sec. 4.3 is $\mathcal{O}(n|\mathcal{C}|^{2}+nt_{\text{LLM-i}})$, (c) The complexity of the two-step sampling in Sec. 4.4 is $\mathcal{O}(\theta n)$. Complete runtime results of LLaTA across all datasets are provided Appendix M.
To practically evaluate the efficiency of LLaTA, we compare its training and inference time with existing all LLM-based GSL methods. The training time includes the fine-tuning process of the LLM and the training of the edge predictor, while the inference time encompasses LLM inference and other GSL modules. As shown in Fig. 5, the experimental results demonstrate that our method requires significantly less inference time, particularly when compared to LLM4RGNN. Furthermore, our approach eliminates the need for any training during the structure learning phase, resulting in zero training time. This highlights the efficiency of our method in seamlessly integrating LLMs with GSL.
@@ Appendix A Encoding Tree and Structural Entropy
(1) The root node $\lambda$ has a partition (community) $\mathcal{P}_{\lambda}^{tree}=\mathcal{V}$, where $\mathcal{V}$ represents the set of all vertices in $\mathcal{G}$.
(2) Each non-root node $\alpha$ has a partition $\mathcal{P}_{\alpha}^{tree}\subset\mathcal{V}$. If $\alpha$ is a leaf node, $\mathcal{P}_{\alpha}^{tree}$ is a singleton with one vertex from $\mathcal{V}$.
(3) For each leaf node $\alpha$ with $\mathcal{P}_{\alpha}^{tree}=v_{i}$, $x_{\alpha}=x_{i}$ is the initial embedding of $\alpha$, and $\mathrm{t}_{\alpha}=\mathrm{t}_{i}$ is the raw text of $\alpha$.
(4) For each non-root node $\alpha$, the parent node of the community to which it belongs is denoted as $\alpha^{+}$.
@@
(6) For each non-leaf node $\alpha$, assuming it has $m$ children, the partitions of its children $\mathcal{P}_{\alpha^{\langle i\rangle}}^{tree}$ together form the partition of $\mathcal{P}_{\alpha}^{tree}$, so that $\mathcal{P}_{\alpha}^{tree}=\bigcup_{i=1}^{m}\mathcal{P}_{\alpha^{\langle i\rangle}}^{tree}$ and $\bigcap_{i=1}^{m}\mathcal{P}_{\alpha^{\langle i\rangle}}^{tree}=\varnothing$. If the height of the encoding tree is restricted to $K$, it is called a $K$-level encoding tree.
@@
where $g_{\alpha}$ represents the sum of the weights of cross-node edges that connect nodes within partition $\mathcal{P}_{\alpha}^{tree}$ to nodes outside $\mathcal{P}_{\alpha}^{tree}$. $\operatorname{vol}(\alpha)$ denotes the sum of the degrees of all nodes within $\mathcal{P}_{\alpha}^{tree}$. $\mathcal{H}^{\mathcal{T}}(\mathcal{G},\alpha)$ is the structural entropy of node $\alpha$, representing the $K$-dimensional structural entropy of $\alpha$ when the height of $\mathcal{T}$ is $K$. The core of this measurement lies in the observation that, in a highly connected graph, nodes frequently interact with their neighbors. By employing random walks, these interactions can be captured, and entropy can be introduced as a measure of topology uncertainty [13]. Specifically, the one-dimensional structural measurement of $\mathcal{G}$ can be quantified using the stationary distribution of its degrees $d$ and Shannon entropy. This concept can be generalized to $K$-dimensional measurements using the structural encoding tree in Definition 1.

### v4

@@ 4.3 Leaf-oriented Two-step Sampling
We perform two-step sampling under the guidance of the following two probability functions:
$P_{topo}(\alpha)=\frac{\mathrm{exp}(\mathcal{H}^{\mathcal{\mathcal{T}}}(\mathcal{G},\alpha))}{{\underset{\gamma\in\mathcal{C}^{\ell}_{\alpha^{+}}}{\sum}}{\mathrm{exp}(H^{\mathcal{\mathcal{T}}}(\mathcal{G},\gamma))}},\;\;P_{sema}^{\alpha}(\beta)=\frac{\exp(\operatorname{sim}(y^{cls}_{\beta},{y}^{cls}_{\alpha}))}{{\underset{\gamma\in\mathcal{C}^{\ell}_{\alpha^{+}},\gamma\neq\alpha}{\sum}}\!\!\!{\mathrm{exp}(\mathrm{sim}(y^{cls}_{\gamma},y^{cls}_{\alpha}))}}.$ (7)
Topology Uncertainty Sampling. The core of GSL is to eliminate structural noise and enhance data utility. Notably, this noise usually appears only in certain substructures rather than the entire graph. According to Theorem 4, nodes with higher structural entropy are more likely to lie in topologically uncertain regions and therefore benefit more from structure refinement. Based on $P_{topo}(\alpha)$ in Eq. 7, we prioritize nodes for sampling within each leaf community $\mathcal{C}^{\ell}$ according to their structural entropy, improving efficiency by focusing on nodes most in need of optimization.
Semantic Similarity Sampling. After selecting a node $\alpha$, we further determine candidate neighbors using $P_{sema}^{\alpha}(\beta)$ in Eq. 7, which measures LLM-empowered semantic similarity between soft labels $y^{cls}$. A $\theta$-sized candidate set is formed from the remaining leaf nodes, and edges are updated accordingly: for edge addition, nodes are ranked by descending similarity; for edge removal, by ascending similarity. This ensures that edge modifications are guided by both semantic consistency and structural refinement.
@@ 5.1 Experiment Setup
We evaluate LLaTA and 14 GSL baselines on 11 widely adopted TAG datasets across multiple domains. The details on these datasets and baselines can be found in Appendix G-H. Due to space limitations, more experimental setup and hyperparameter details are provided in Appendix I.
@@ 5.2 Performance Comparison
Table 1: Node classification accuracy(%) on 11 TAG datasets. Highlighted are the top first, second, and third accuracy. "OOM" denotes out of memory and "OOT" denotes out of time.
Method Cora Citeseer Pubmed WikiCS Instagram Reddit Ratings Child History Photo ArXiv IDGL 86.81±0.31 75.39±0.27 87.25±0.22 79.78±0.14 63.68±0.09 65.03±0.13 43.29±0.29 45.33±0.26 80.46±0.31 82.33±0.33 OOM SLAPS 79.89±0.50 73.51±0.64 86.41±0.49 72.50±0.27 60.91±0.13 OOM 40.76±0.23 47.46±0.21 OOM OOM OOM GAUGO 85.24±0.27 75.24±0.43 87.12±0.38 70.74±0.29 63.34±0.21 OOM 40.19±0.24 47.27±0.19 OOM OOM OOM HESGSL 85.91±0.45 76.03±0.31 87.17±0.27 73.29±0.22 63.83±0.16 OOM 38.71±0.42 47.81±0.25 OOM OOM OOM SEGSL 86.90±0.54 76.52±0.20 87.42±0.38 79.68±0.29 65.26±0.21 64.87±0.26 43.20±0.47 51.80±0.39 OOT OOT OOT ProGNN 84.18±0.23 75.08±0.21 86.89±0.14 71.90±0.16 64.45±0.19 OOM OOM OOM OOM OOM OOM SUBLIME 84.81±0.33 75.55±0.47 87.63±0.70 78.06±0.38 64.78±0.24 58.82±0.21 39.90±0.17 49.54±0.16 OOM OOM OOM STABLE 84.21±0.43 75.34±0.60 87.27±0.39 76.93±0.33 65.66±0.18 59.78±0.24 40.55±0.10 50.90±0.09 OOM OOM OOM CoGSL 86.16±0.45 75.45±0.36 86.52±0.49 OOM OOM OOM OOM OOM OOM OOM OOM BORF 87.08±0.18 76.43±0.17 87.45±0.11 80.56±0.08 64.07±0.06 63.07±0.09 42.72±0.60 50.92±0.54 82.83±0.14 82.63±0.15 73.08±0.33 DHGR 86.72±0.42 76.18±0.14 86.76±0.33 79.25±0.17 64.26±0.26 66.41±0.29 43.33±0.36 52.41±0.40 82.61±0.16 81.92±0.18 73.47±0.28 GraphEdit 86.26±1.06 76.33±1.12 87.14±0.29 79.92±0.63 64.23±1.01 OOT 42.53±0.97 OOT OOT OOT OOT LLM4RGNN 86.73±0.73 76.96±0.58 87.37±0.21 OOT 64.85±0.66 OOT OOT OOT OOT OOT OOT LangGSL 87.65±0.67 76.69±0.73 87.24±1.21 80.26±0.45 66.25±0.39 66.87±0.12 44.01±1.37 49.66±1.05 OOT OOT OOT LLaTA (Ours) 88.26±0.26 78.21±0.18 88.39±0.23 81.58±0.20 66.73±0.13 67.60±0.19 44.47±0.08 62.28±0.56 85.28±0.24 85.41±0.24 75.39±0.41
@@
Method Cora Citeseer Pubmed Instagram SEGSL 67.70±0.17 71.15±0.12 79.05±0.04 64.63±0.01 BORF 67.54±0.16 70.62±0.10 79.46±0.03 64.49±0.02 DHGR 66.85±0.63 70.14±0.21 78.64±0.15 64.52±0.06 LLM4RGNN 67.63±0.11 71.72±0.13 78.90±0.06 OOT LangGSL 67.82±0.16 71.38±0.14 79.95±0.07 65.03±0.01 LLaTA 68.39±0.14 72.18±0.15 80.25±0.06 65.37±0.02
@@ 5.4 Robustness Analysis
Downstream Backbones. According to Table 5, LLaTA consistently improves performance of the backbones across all four datasets, with gains ranging from 0.56% to 10.18%. These enhancements are observed not only on conventional GNN backbones (GCN, GAT, GraphSAGE) but also on LLM-GNN backbones (GLEM, ENGINE), highlighting the effectiveness and applicability of our structure optimization across diverse downstream architectures.
Table 4: Improvements of LLaTA over different GNN/LLM-GNN backbones.
[…]
@@
Method Time Complexity SEGSL $\mathcal{O}(n^{2}+nlog^{2}n+n)$ GraphEdit $\mathcal{O}(pt_{\text{LLM-r}}+qd^{2}+n^{2}+(m+nk)t_{\text{LLM-i}})\;\;\;\;\;\;$ LLM4RGNN $\mathcal{O}(pt_{\text{LLM-r}}+mt_{\text{LLM-i}}+qd^{2}+n^{2})$ LangGSL $\mathcal{O}(nt_{\text{LLM-i}}+nt_{\text{LM-e}}+n^{2}+nd^{2}+n^{2}d)$ LLaTA $\mathcal{O}(nlog^{2}n+n|\mathcal{C}|^{2}+nt_{\text{LLM-i}}+n)$
Figure 3: Node classification performance on the Cora and Pubmed dataset under real-world scenarios (Sparsity [Edge Removal] and Noise [Edge Addition]).
Real-world Scenarios. To evaluate the robustness of LLaTA against sparsity and noise scenarios in practical applications, we randomly remove or add edges to the original graph structure. As illustrated in Fig. 3, LLaTA demonstrates strong predictive performance in both scenarios. In the sparse setting, LLaTA achieves optimal performance across most perturbation levels. In the noisy scenario, LLaTA performs competitively, with only LLM4RGNN surpassing it, as the latter is specifically designed to handle graph adversarial attack settings. Notably, as the edge addition rate increases, the performance of LLaTA gradually deteriorates, which can be attributed to the irreversible negative impact of such perturbations on the tree initialization. Further experimental results are provided in the Appendix L.
Hyperparameter. We conducted a comprehensive analysis of the four key hyperparameters of LLaTA: $K,\epsilon,\theta$, and $r$. The experimental results for $K,\theta$, and $r$ are presented in Fig. 5, while the more experimental analysis and the detailed selection guidance of the four hyperparameters are provided in Appendix M. In Fig. 5 (left), the optimal tree height $K$ differs across datasets, reflecting varying graph complexity. However, when $K$ reaches 6, performance drops markedly for all datasets, suggesting that excessive tree depth may cause overfitting or loss of structural information. In Fig. 5 (right), we analyze LLaTA’s sensitivity to the candidate set size $\theta$ in semantic similarity sampling and the two-step sampling frequency $r$ on the History dataset. The results indicate that an $r$ value between 3 and 10 yields higher accuracy, corresponding to an appropriately improved graph density. In contrast, LLaTA is less sensitive to the choice of $\theta$, achieving stable performance within the range of $\theta\in[5,15]$. A smaller $\theta$ may result in insufficient semantic similarity sampling, whereas a larger $\theta$ could introduce noise into the candidate set, negatively impacting performance.
Figure 4: Hyperparameter analysis of $K$, $\theta$ and $r$. $\theta$ and $r$ are analyzed on the History dataset.
@@ 5.5 Efficiency analysis
In Table 5, we analyze the theoretical time complexity of LLaTA and existing LLM-based GSL methods. Where $q$ and $p$ is the number of samples for fine-tuning and edge predictor training, $d$ denotes the feature dimension and $t_{\text{LM-e}}$ represents the LM encoding time. $t_{\text{LLM-r}}$ and $t_{\text{LLM-i}}$ denote the fine-tuning and inference time of LLM, respectively. The magnitude of $t_{\text{LLM-i}}$ should be adaptively determined according to the specific requirements of different inference tasks. For LLaTA, The time complexity can be divided into three parts: (a) The complexity for tree construction in Sec. 4.1 is $\mathcal{O}(nlog^{2}n)$, (b) The complexity of tree-prompted LLM inference in Sec. 4.2 is $\mathcal{O}(n|\mathcal{C}|^{2}+nt_{\text{LLM-i}})$, (c) The complexity of the two-step sampling in Sec. 4.3 is $\mathcal{O}(\theta n)$. Complete runtime results and scalability analysis of LLaTA across all datasets are provided Appendix N.
To practically evaluate the efficiency of LLaTA, we compare its training and inference time with existing LLM-based GSL methods. The training time includes the fine-tuning process of the LLM and the training of the edge predictor, while the inference time encompasses LLM inference and other GSL modules. As shown in Fig. 5, the experimental results demonstrate that our method requires significantly less inference time, particularly when compared to LLM4RGNN. Furthermore, our approach eliminates the need for any training during the structure learning phase, resulting in zero training time. This highlights the efficiency of our method in seamlessly integrating LLMs with GSL.
@@ Appendix A Encoding Tree and Structural Entropy
(1) The root node $\lambda$ has a community $\mathcal{C}_{\lambda}=\mathcal{V}$, where $\mathcal{V}$ represents the set of all vertices in $\mathcal{G}$.
(2) Each non-root node $\alpha$ corresponds to a community $\mathcal{C}_{\alpha}\subset\mathcal{V}$. If $\alpha$ is a leaf node, $\mathcal{C}_{\alpha}$ is a singleton with one vertex from $\mathcal{V}$.
(3) For each leaf node $\alpha$ with $\mathcal{C}_{\alpha}=v_{i}$, $x_{\alpha}=x_{i}$ is the initial embedding of $\alpha$, and $\mathrm{t}_{\alpha}=\mathrm{t}_{i}$ is the raw text of $\alpha$.
(4) For each non-root node $\alpha$, the parent node corresponding to its community which it belongs to is denoted as $\alpha^{+}$.
@@
(6) For each non-leaf node $\alpha$, assuming it has $m$ children, the communities of its children $\mathcal{C}_{\alpha^{\langle i\rangle}}$ together form the community of $\mathcal{C}_{\alpha}$, so that $\mathcal{C}_{\alpha}=\bigcup_{i=1}^{m}\mathcal{C}_{\alpha^{\langle i\rangle}}$ and $\bigcap_{i=1}^{m}\mathcal{C}_{\alpha^{\langle i\rangle}}=\varnothing$. If the height of the encoding tree is restricted to $K$, it is called a $K$-level encoding tree.
@@
where $g_{\alpha}$ represents the sum of the weights of cross-node edges that connect nodes within partition $\mathcal{C}_{\alpha}$ to nodes outside $\mathcal{C}_{\alpha}$. $\operatorname{vol}(\alpha)$ denotes the sum of the degrees of all nodes within $\mathcal{C}_{\alpha}$. $\mathcal{H}^{\mathcal{T}}(\mathcal{G},\alpha)$ is the structural entropy of node $\alpha$, representing the $K$-dimensional structural entropy of $\alpha$ when the height of $\mathcal{T}$ is $K$. The core of this measurement lies in the observation that, in a highly connected graph, nodes frequently interact with their neighbors. By employing random walks, these interactions can be captured, and entropy can be introduced as a measure of topology uncertainty [12]. Specifically, the one-dimensional structural measurement of $\mathcal{G}$ can be quantified using the stationary distribution of its degrees $d$ and Shannon entropy. This concept can be generalized to $K$-dimensional measurements using the structural encoding tree in Definition 1.